# CRUD-RAG: A Comprehensive Chinese Benchmark for Retrieval-Augmented Generation of Large Language Models

### v2

@@ 4.5. Analyzing the Impact of Embedding Model on RAG Performance in Different Tasks
Most RAG systems use vector similarity-based algorithms as retrievers. Therefore, the embedding model that converts document blocks into vectors is crucial for the retrieval effect. We tested various embedding models optimized for retrieval tasks and compared their performance in the RAG system. We evaluated several embedding models with similar parameter sizes. According to  (Muennighoff et al., 2022), the embedding models’ performance on the retrieval task should follow the order of gte ¿ stella ¿ bge ¿ m3e. Our benchmark confirmed this order, but with some variations.
@@
However, for question answering tasks that required precise localization of relevant documents, we found that m3e-base performed much worse than gte-base. This matched the finding of  (Muennighoff et al., 2022).
More importantly, for the hallucination modification task, m3e-base, which ranked the lowest on the retrieval benchmark, outperformed the other models on all metrics. This implies that m3e-base is better at retrieving relevant content for tasks that correct entity-level errors in the text.
These results further show that the retrieval benchmark may not be fully appropriate for RAG. In general, question answering tasks that need accurate localization of relevant documents can use the retrieval benchmark to select models.
@@ 4.6. Analyzing the Impact of Top-k on RAG Performance in Different Tasks
The RAG system converts the user’s query into a vector using the same embedding model as the vector database. Then, it searches the index for the top-k most similar vectors to the query vector , and retrieves the corresponding text blocks from the database. These text blocks serve as the context for the LLM prompt. The amount of information that the model receives depends on the size of k. We will show how the amount of context information affects the system performance for different tasks in the Table 7.
Text Continuation: Text continuation is a highly creative task. Table 7 shows that increasing top-k improves both the overall semantic similarity metrics (bertScore, bleu and rouge-L) and the RAGQuestEval metrics. The recall metric of RAGQuestEval shows how much key information from the reference is included in the generated text, while the precision metric shows how correct and relevant that information is. We found that higher top-k values lead to higher recall and precision scores, indicating that the generated text contains more and better key information. We attribute this to the increased diversity and accuracy of the generated text from more documents.
@@
Question Answering: For single-document QA, increasing top-k has little impact on the semantic similarity metric, but improves the RAGQuestEval metrics, which measure the accuracy and recall of key information. Higher top-k values lead to higher recall and precision scores, especially when the retrieved content is small. This is because more documents can confirm the answer multiple times.
@@ 4.7. Analyzing the Impact of LLM on RAG Performance in Different Tasks
The core of the RAG system is a large language model (LLM), which can generate accurate and fluent answers based on the user’s question and the retrieved information. In this paper, we conducted experiments on several commonly used LLMs and showed their performance on our benchmark.
@@
Question Answering: For single-document QA, which only require extracting relevant information from a sentence in the text, this task is relatively simple. Qwen and Baichuan2 even outperformed GPT series models. However, for multi-document QA that require comprehensive understanding of multiple documents, GPT-4 was far ahead of other models, showing its excellent knowledge fusion ability. Baichuan2-13B model also performed better than GPT-3.5, indicating its potential.
Hallucination Modification: We found that some models generated text that was too long, introducing redundant information. The hallucination modification task only requires modifying the hallucination information, retaining other information, and not introducing irrelevant information. Therefore, ChatGLM2, Qwen-7B, Baichuan2 did not complete this task well.
In summary, the GPT-4-0613 model performed excellently on most tasks and evaluation metrics, proving that it is a powerful LLM. Qwen-7B and Qwen-14B models also performed well, especially in the text continuation and summary generation tasks. Baichuan2-13B model was very competitive with GPT-4 in the QA task, deserving more investigation.
4.8. Suggestions on how to optimize your RAG system
@@
Create: Creative Content Generation
Larger, to access diverse knowledge
Larger, to preserve article structure
Larger, to maintain semantic coherence
Dense algorithm for semantic understanding
Qwen-14B for cost-effective high-quality text
Delete: Summarization
Moderate, for precision-recall balance
Smaller for more recall, larger for more precision
Larger, to maintain semantic coherence
BM25 for precise content, Dense algorithm for more recall
Qwen-14B for high-quality summaries
Read: Single-document QA
Larger, for repeated determination
Moderate, for pinpointing short answers
Larger, to maintain semantic coherence
Hybrid + rerank for enhanced performance
Baichuan2-13B for GPT-4-like performance
Read: Multi-document QA
Larger, for retrieving complementary articles
Larger, for article completeness
Larger, to maintain semantic coherence
Hybrid + rerank for enhanced performance
Baichuan2-13B for GPT-4-like performance
Update: Error Correction
Smaller, for high precision tasks
Larger, to avoid breaking article structure
Smaller, error correction tasks are not sensitive to semantic coherence
BM25 for precise content generation
GPT-4 or alternatives depending on cost
The chunk size is also an important factor when building the vector index for external knowledge. For creative scenarios, such as content generation, we suggest using a larger chunk size to preserve the structure of the article and avoid affecting the performance of the RAG system. For summary scenarios, a smaller chunk size can be used if more information is desired to be recalled; however, if the precision of the generated content is more important, a larger chunk size is still recommended to avoid destroying the structure of the article. In factual question answering scenarios, a smaller chunk size is beneficial for finding the answer in a short sentence. For reasoning tasks, a larger chunk size can ensure the article completeness and enhance the reasoning ability.
@@
Moreover, the hybrid algorithm that combines dense and BM25 retriever has very limited improvement on the overall quality of the generated results. However, by using a rerank model to reorder the retrieval results and then inputting them into LLMs, the performance of almost all tasks improved, especially for reasoning tasks. Therefore, we suggest trying to use the hybrid algorithm + rerank retrieval mode when the conditions permit, which can achieve better performance in the RAG system.
When choosing a large language model, GPT-4 model is undoubtedly the most advanced model at present. However, due to the high cost of invoking GPT-4, we may need to consider some open-source alternatives. According to our experimental results, Qwen-14B model has shown similar performance to GPT-4 in the two tasks of text continuation and summary generation, and can generate high-quality creative and summarizing texts. And in the QA task, Baichuan2-13B model also showed a level close to GPT-4, and can generate accurate and fluent answers. Therefore, we can choose the suitable large language model according to different tasks and cost requirements.

### v3

@@ 4.4. Analyzing the Impact of Retriever on RAG Performance in Different Tasks
Text Continuation: As Table 5 displays, the performance of the dense retriever is roughly equivalent to that of BM25, except for the key information recall rate. Compared to the keyword-based algorithm, the modern vector search can capture the semantic and contextual information of words, so that more content that does not match keywords but is obviously semantically related can be retrieved. However, the RAG system using BM25 also performs well. In terms of the precision of key information, BM25 even exceeds the dense retriever. This suggests that in the continuation task, which is a creative task, BM25 can retrieve content that is highly relevant to the user’s intention, but may overlook some details.
Open-Domain Multi-Document Summarization: On the overall semantic similarity metric, the performance of the dense retriever is roughly equivalent to that of BM25. On the QuestEval metric, BM25 surpasses dense retriever in terms of key information precision, but slightly trails behind in key information recall. If the retrieved content contains a lot of irrelevant information, the model-generated summary may have errors or redundancies. BM25 retrieved content usually matches the user’s intention better, but sometimes may miss some important information. Therefore, BM25 is weaker than dense retriever in key information recall, but excels in key information accuracy. Besides, hybrid retrieval algorithms presumably combine the advantages of both, and the RAG system generates content with suitable precision and recall.
Question Answering: In question answering, we find that dense retriever has a more obvious advantage over BM25, when dealing with reasoning questions that require synthesizing multiple documents. In question-answering tasks that require considering three documents, Dense retriever not only surpasses BM25 in all the overall semantic similarity metrics, but also achieves a significant improvement in key information precision and recall. This indicates that question-answering retrieval is more difficult than text continuation and other tasks, especially reasoning question-answering, which requires a higher level of semantic understanding, and simple keyword retrieval algorithms may not be sufficient. We also found that the Hybrid+Rerank algorithm, which combines and re-ranks the results of both algorithms, improves on all evaluation metrics. This suggests that this is a better retrieval algorithm for question-answering tasks.
Hallucination Modification: Consistent with the conclusion of summarization, the BM25 retriever performs slightly better than or equal to the dense retriever. For RAG tasks such as hallucination modification, which require precise retrieval of highly relevant content, BM25 shows good performance. Moreover, BM25 requires less computational resources than dense retrievers. This indicates that different RAG tasks require different retrieval algorithms.
Retrieval Accuracy Evaluation: To make a more comprehensive evaluation, we evaluated the retrieval accuracy on question answering and hallucination modification tasks using MRR (mean reciprocal rank) as a separate metric. This separate evaluation allows for a more accurate assessment of the retriever’s capabilities. Notably, text continuation and open-domain summarization tasks were excluded due to their subjective and vague evaluation criteria, lacking clear ground truth. Additionally, both 2-document and 3-document question answering require multiple documents to address queries. Therefore, we calculate the MRR for each retrieved document individually and take the average as the final result. The pure hybrid algorithm was not evaluated separately as it could alter the order of retrieved content, affecting subsequent processing steps.
Figure 7 shows that the hybrid + reranking method excels in most tasks, outperforming other methods. This demonstrates the effectiveness of combining multiple retrieval strategies with reranking. Notably, BM25 and dense retrievers perform comparably in many cases, highlighting the strengths of both traditional and neural network methods. In question answering, performance for all methods declines as the number of documents increases, aligning with expectations since multi-document tasks are more challenging and require stronger information integration. These results are consistent with our previous end-to-end evaluations, confirming the reliability of the end-to-end evaluation method.
Figure 8. Comparison of Mean Reciprocal Rank (MRR) scores for different embedding models in our benchmark.
@@ 4.5. Analyzing the Impact of Embedding Model on RAG Performance in Different Tasks
Most RAG systems use vector similarity-based algorithms as retrievers. Therefore, the embedding model that converts document blocks into vectors is crucial for the retrieval effect. We tested various embedding models optimized for retrieval tasks and compared their performance in the RAG system. We evaluated several embedding models with similar parameter sizes. According to  (Muennighoff et al., 2022), the embedding models’ performance on the retrieval task should follow the order of GTE ¿ STELLA ¿ BGE ¿ M3E. Our results show some variations with this order.
@@
For single-document question answering tasks that required precise localization of relevant documents, we found that m3e-base performed much worse than others. This matched the finding of  (Muennighoff et al., 2022). However, for the hallucination modification task, m3e-base, which ranked the lowest on the retrieval benchmark, outperformed the other models on all metrics. These results further show that the retrieval benchmark may not be fully appropriate for RAG.
Retrieval Accuracy Evaluation: Similar to the experiments in the retriever evaluation, we use the MRR metric to evaluate four mainstream embedding methods: BGE, GTE, M3E, and STELLA. The results in Figure 8 indicate that the performance of these methods is relatively close across different tasks, with no single method outperforming the others in all tasks. This underscores the importance of considering specific task requirements when selecting an embedding method.
As the number of documents increases from 1 to 3, the MRR values for all methods show a downward trend. This trend aligns with our previous end-to-end experimental results, highlighting the challenges of multi-document understanding tasks.
@@ 4.6. Analyzing the Impact of Top-k on RAG Performance in Different Tasks
The RAG system converts the user’s query into a vector using the same embedding model as the vector database. Then, it searches the index for the top-k most similar vectors to the query vector, and retrieves the corresponding text blocks from the database. These text blocks serve as the context for the LLM prompt. The amount of information that the model receives depends on the size of k. We will show how the amount of context information affects the system performance for different tasks in Table 7.
Text Continuation: Text continuation is a highly creative task. Table 7 shows that increasing top-k improves both the overall semantic similarity metrics (bertScore, bleu, and rouge-L) and the RAGQuestEval metrics. The recall metric of RAGQuestEval shows how much key information from the reference is included in the generated text, while the precision metric shows how correct and relevant that information is. We found that higher top-k values lead to higher recall and precision scores, indicating that the generated text contains more and better key information. We attribute this to the increased diversity and accuracy of the generated text from more documents.
@@
Question Answering: For single-document QA, increasing top-k has little impact on the semantic similarity metric, but improves the RAGQuestEval metrics, which measure the accuracy and recall of key information. When the top-k value is too small, increasing the value of the top-k can significantly increase the recall and precision scores. This is because when the retrieved content is small, it may not be helpful for answering.
@@ 4.7. Analyzing the Impact of LLM on RAG Performance in Different Tasks
The core of the RAG system is an LLM, which can generate accurate and fluent answers based on the user’s question and the retrieved information. In this paper, we conducted experiments on several commonly used LLMs, as Table  8 displayed.
@@
Question Answering: For single-document QA, which only requires extracting relevant information from a sentence in the text, this task is relatively simple. Qwen and Baichuan2 even outperformed GPT series models. However, for multi-document QA that requires a comprehensive understanding of multiple documents, GPT-4 was far ahead of other models, showing its excellent knowledge fusion ability. The Baichuan2-13B model also performed better than GPT-3.5, indicating its potential.
Hallucination Modification: We found that some models generated text that was too long, introducing redundant information. The hallucination modification task only requires modifying the hallucination information, retaining other information, and not introducing irrelevant information. Therefore, ChatGLM2, Qwen-7B, and Baichuan2 did not complete this task well.
In summary, the GPT-4 model performed excellently on most tasks and evaluation metrics, proving that it is a powerful LLM. Qwen-7B and Qwen-14B models also performed well, especially in the text continuation and summary generation tasks. Baichuan2-13B model was very competitive with GPT-4 in the QA task, deserving more investigation.
Latest LLM Evaluation: Our dataset was constructed in December 2023. To evaluate its challenge to the latest LLMs in 2024, we experimented with two newly released models: GPT-4o(Released in May 2024) and Qwen2-7b(Released in June 2024).
The results show that GPT-4o performs similarly to its predecessor GPT-4, or with some slight improvements. In contrast, Qwen2-7b demonstrates significant improvements over its predecessor Qwen-7b in multiple tasks. These findings confirm that our benchmarks remain challenging for the latest LLMs. Additionally, it is encouraging to observe that the performance of many LLMs continues to improve with each new version.
4.8. Suggestions for Optimizing Your RAG System
@@
Create: Creative Content Generation Larger, to access diverse knowledge Larger, to preserve article structure Larger, to maintain semantic coherence Dense algorithm for semantic understanding Qwen-14B for cost-effective high-quality text
Delete: Summarization Moderate, for precision-recall balance Smaller for more recall, larger for more precision Larger, to maintain semantic coherence BM25 for precise content, Dense algorithm for more recall Qwen-14B for high-quality summaries
Read: Single-document QA Larger, for repeated determination Moderate, for pinpointing short answers Larger, to maintain semantic coherence Hybrid + rerank for enhanced performance Baichuan2-13B for GPT-4-like performance
Read: Multi-document QA Larger, for retrieving complementary articles Larger, for article completeness Larger, to maintain semantic coherence Hybrid + rerank for enhanced performance Baichuan2-13B for GPT-4-like performance
Update: Error Correction Smaller, for high precision tasks Larger, to avoid breaking article structure Smaller, error correction tasks are not sensitive to semantic coherence BM25 for precise content generation GPT-4 or alternatives depending on cost
The chunk size is also an important factor when building the vector index for external knowledge. For creative scenarios, such as content generation, we suggest using a larger chunk size to preserve the structure of the article and avoid affecting the performance of the RAG system. For summary scenarios, a smaller chunk size can be used if more information is desired to be recalled; however, if the precision of the generated content is more important, a larger chunk size is still recommended to avoid destroying the structure of the article. In factual question answering scenarios, a smaller chunk size is beneficial for finding the answer in a short sentence. For reasoning tasks, a larger chunk size can ensure the article’s completeness and enhance the reasoning ability.
@@
Moreover, the hybrid algorithm that combines dense and BM25 retriever has very limited improvement on the overall quality of the generated results. However, by using a rerank model to reorder the retrieval results and then inputting them into LLMs, the performance of almost all tasks improved, especially reasoning tasks. Therefore, we suggest trying to use the hybrid algorithm + rerank retrieval mode when the conditions permit, which can achieve better performance in the RAG system.
When choosing a large language model, GPT-4 model is undoubtedly the most advanced model at present. However, due to the high cost of invoking GPT-4, we may need to consider some open-source alternatives. According to our experimental results, Qwen-14B model has shown similar performance to GPT-4 in the two tasks of text continuation and summary generation, and can generate high-quality creative and summarizing texts. In the QA task, Baichuan2-13B model also showed a level close to GPT-4, and can generate accurate and fluent answers. Therefore, we can choose a suitable LLM according to different tasks and cost requirements.